# Generative diffusion posterior sampling for informative likelihoods

### v1

@@ Abstract
Generative diffusion models, Feynman–Kac models, conditional sampling, posterior sampling, sequential Monte Carlo,
@@ 1 Problem formulation
The overall goal of the paper is to sample from the conditional/posterior distribution $\pi(\cdot{\;|\;}y)\propto f(y{\;|\;}\cdot)\,\pi(\cdot)$, where $f$ is a given likelihood function, and $y\in\mathbb{R}^{c}$ denotes an observation (Zhao et al.,, 2025). This task is at the core interests of Bayesian statistics in general, and in the context of generative diffusion models specifically, this has enabled many downstream applications, such as image restoration problems (Luo et al.,, 2025, 2023). Recently, a few methods have been proposed to address this problem, notably the diffusion posterior sampling (Chung et al.,, 2023) and their variants (see survey by Daras et al.,, 2024). These methods rely on constructing a conditional diffusion process characterised by transition $\widetilde{q}_{k{\;|\;}k-1}(u_{k}{\;|\;}u_{k-1},y)$ that encodes the likelihood $f$ and observation, to target $\pi(\cdot{\;|\;}y)$. However, these methods are intrinsically biased due to necessary approximations to $\widetilde{q}_{k{\;|\;}k-1}$. This has then motivated another class of methods based on Feynman–Kac models and sequential Monte Carlo (SMC, Chopin and Papaspiliopoulos,, 2020), to remove such biases (e.g., Wu et al.,, 2023; Cardoso et al.,, 2024; Janati et al.,, 2024; Kelvinius et al.,, 2025; Corenflos et al.,, 2025). At the heart, they work by using $\widetilde{q}$ as a proposal, and then correct the proposal samples with importance weights. While this type of methods has shown to be empirically working well, their statistical efficiencies (e.g., effective sample size) can be significantly degraded by high dimension $d$ and informative likelihood. This challenge has long been recognised in the SMC literature, such as the seminal work by Del Moral and Murray, (2015). However, in the context of SMC conditional sampling of generative diffusion models, this is under investigated. As such, the specific purpose of this paper is to address the problem when the likelihood being highly informative or when the observation is an outlier.
@@
The paper is organised as follows. In Section 2 we introduce necessary premises and explain how the conditional sampling works with Feynman–Kac models and SMC samplers. Then, in Section 3 we present our construction of the Feynman–Kac model to facilitate sampling under outlier observations, followed by experiments in Section 4.
@@ 2 Generative Feynman–Kac models
where $M_{0}$ is the initial distribution, $Z_{N}$ is the normalising constant, and $M_{k{\;|\;}k-1}$ and $G_{k}$ are Markov (proposal) transition distribution and potential function, respectively. Given the model components, a natural approach for sampling $Q_{0:N}$ is sequential Monte Carlo (SMC) which exploits the model’s temporal structure, see Algorithm 1 for implementation. At its heart, this can be viewed as a tempering sequence, gradually annealing from $Q_{0}$ to $Q_{N}$. For detailed expositions of Feynman–Kac models and SMC samplers, we refer the readers to Chopin and Papaspiliopoulos, (2020) and Del Moral, (2004).
@@
A canonical, and perhaps also the most used twisting function (Wu et al.,, 2023) is given by
@@
such that the marginal distribution $Q_{k}(u_{k})=\pi(u_{k}{\;|\;}y)$ represents the conditional denoising model. Notably, The resulting Feynman–Kac model gives an Eulerian representation of the conditional score stochastic differential equation (SDE, Song et al.,, 2021, Equ. 14) at discrete times when the denoising model $q_{k{\;|\;}k-1}$ is given by a score SDE. Precisely, if the noising process is an SDE $\mathop{}\!\mathrm{d}X_{t}=a(X_{t},t)\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}B_{t}$ with $X_{0}\sim\pi$, then $Q_{0:N}$ is the finite-dimensional distribution of the conditional SDE
$\begin{split}\mathop{}\!\mathrm{d}U_{t}&=-a(U_{t},T-t)+\bigl{(}\operatorname{\nabla\!}\log p_{T-t}(U_{t})+\operatorname{\nabla\!}\log l^{y}_{T-t}(U_{t})\bigr{)}\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}W_{t},\\ U_{t}&\sim\pi_{\mathrm{ref}}(\cdot{\;|\;}y),\end{split}$ (8)
at discrete times $t_{0},t_{1},\ldots,t_{N}$. The score $\operatorname{\nabla\!}\log l^{y}_{k}(\cdot)$, sometimes referred to as a guidance or control term, is added alongside with the unconditional score function $\operatorname{\nabla\!}\log p_{k}(\cdot)$ in the conditional SDE. Although the exact twisting function in Equation (7) is intractable in practice, many useful and computational approximations have been established (see surveys by Daras et al.,, 2024; Luo et al.,, 2025). Furthermore, these approximations can be directly integrated in the Feynman–Kac model, for improved statistical performance (e.g., Kelvinius et al.,, 2025) with SMC samplers. These are the reasons why this canonical twisting construction is popularly used. As an example, Wu et al., (2023) combine Tweedie’s formula and first-order approximation to the integral: $l^{y}_{k}(u_{k})\approx f(y{\;|\;}\operatorname{\mathbb{E}}[X_{0}{\;|\;}X_{N-k}=u_{k}])$ which essentially generalises that of Chung et al., (2023).
However, many criticism can be said to this canonical twisting. First, in almost all practical cases, the canonical twisting has to be approximated, and the approximation errors are substantial and difficult to control particularly when $N-k$ is large. Second, even if we were able to compute the canonical twisting exactly, constructing an effective proposal $M_{k{\;|\;}k-1}$, see Remark 2.1, is hard as well. These two problems are especially pronounced when the likelihood $f$ is informative, or $y$ is an outlier observation. To build some intuition, consider a heuristic: suppose that we have $J$ samples $\{U^{j}_{0}\}_{j=1}^{J}\sim q_{0}$, then, to get weighed samples for $Q_{0}(\cdot)\propto l^{y}_{0}(\cdot)\,q_{0}(\cdot)$, Equation (7) is essentially pushforwarding $\{U^{j}_{0}\}_{j=1}^{J}$ to $\{U^{j}_{N}\}_{j=1}^{J}$ and then obtain the weights by evaluating $f(y{\;|\;}\{U^{j}_{N}\}_{j=1}^{J})$ which are likely to be degenerate under outlier $y$.
@@ 3 Construction of twisting sequence with observation path
After all, if our primary interest is only the marginal $q_{N}(\cdot{\;|\;}y)$, we are not obliged to constrain the path $Q_{0:N}$ to follow the SDE representation in Equation (8). In this section, we depart from the canonical design, and develop a new construction of the twisting functions $\{l^{v_{k}}_{k}\}_{k=0}^{N}$, especially tailored for informative likelihood or outlier observation. Our gist consists in designing $l^{v_{k}}_{k}$ in a way that it is an “appropriate” likelihood for $q_{k}$ at $k$, by making a smooth bridging/tempering between $l_{0}^{v_{k}}$, the reference likelihood, and $l_{N}^{v_{N}}$, the target likelihood. To clarify, for instance, if $q_{0}$ is a standard Gaussian, then we ideally would like $l_{0}^{v_{0}}\approx\mathrm{N}(0;0,I)$ facilitating importance sampling at this step. We achieve this by drawing inspiration from Corenflos et al., (2025) and Dou and Song, (2024), to construct an auxiliary path of observations correlated with the diffusion model. The intermediate likelihood will then reflect this correlation at all steps for SMC samplers to leverage.
We define a new process $\{Y_{k}\}_{k=0}^{N}$ initialised with the given observation $Y_{0}=y\equiv y_{0}$, and simulate the path $\{Y_{k}\}_{k=0}^{N}$ by a noising transition $p_{k{\;|\;}k-1}(y_{k}{\;|\;}y_{k-1})$. [^1] For simplicity, we let the noising notation for $y$ be the same as in Equation (1) for $x$, but they do not need to be the same in practice. Write its time-reversal as $\{V_{k}=Y_{N-k}\}_{k=0}^{N}$. We design the twisting function as a time-pushforward of the target likelihood:
@@ 3.1 Construction with Gaussian likelihood models
which is in general intractable mainly due to the denoising $p_{0}(x_{0}{\;|\;}x_{k})$. It seems that the same problem of the canonical twisting in Equation (7) persists, as existing approximations to $p_{0}(x_{0}{\;|\;}x_{k})$ will result in large errors due to the time leap from $k$ to 00. However, the advantage of this twisting is that we can alternatively obtain $p_{k}(y_{k}{\;|\;}x_{k})$ based on $p_{k-1}(y_{k-1}{\;|\;}x_{k-1})$ for which we can then recursively compute $p_{k}(y_{k}{\;|\;}x_{k})$ starting at the target likelihood $p_{0}(y_{0}{\;|\;}x_{0})=f(y_{0}{\;|\;}x_{0})$. This is shown in the following proposition.
@@
If the observation noising process $p_{k-1}(y_{k}{\;|\;}y_{k-1})$ is $\phi$-stationary: $\int p_{k-1}(y_{k}{\;|\;}y_{k-1})\mathop{}\!\mathrm{d}\phi(y_{k-1})=\phi(y_{k})$, then $\mathcal{K}(\phi)(y,\cdot)=\phi(y)$ is a fixed point. The resulting twisting function will be approximately a standard Normal for large $N$ if the noising is an OU process.
Recall that the canonical twisting in Equation (7) requires global approximations to $p_{0}(x_{0}{\;|\;}x_{k})$ between 00 and $k$, which is hard. In contrast, the interpolating twisting in Equation (11) only needs local approximations between continuum $k-1$ and $k$, significantly reducing complexity and improving numerical robustness in practice. We show a particular case when the target likelihood is linear Gaussian for how to approximate $\mathcal{K}^{k}(f)$ in closed form.
@@
From now on, we will adopt the proposed twisting function
@@
using the reverse-time notation. Although this twisting is approximate, the marginal $q_{N}(\cdot{\;|\;}y)$ that we are primarily concerned is left un-approximated. Furthermore, this twisting function gives us a tractable guided proposal $M_{k{\;|\;}k-1}$ mentioned in Remark 2.1 allowing for efficiently sampling further:
@@ 2 Reverse $v_{k}=y_{N-k}$ for $k=0,1,\ldots,N$
Compute $\{F_{k},z_{k},\Omega_{k}\}_{k=1}^{N}$ in Proposition 3.2
  // Can do online in $k$
3 Using $M_{k{\;|\;}k-1}$ and $G_{k}$ in Equation (14) and (15)
@@ 4 Experiments
In this section we validate the proposed B0SMC sampler in Algorithm 2 on a high-dimensional ($d=256$) conditional sampling problem. Importantly, we focus on gauging how statistically robust the method is against different levels of outlier observation and number of particles. The implementation is performed on an NVIDIA A100 80G GPU using JAX (Bradbury et al.,, 2018), and the code is publicly available at https://github.com/zgbkdlm/gfk for reproducibility.
The target posterior distribution is defined by a Gaussian mixture model
@@
For more in-depth comparison, we also plot the ESS and time-evolution of marginal $Q_{k}$ in Figures 1 and 2. The first figure gives us three observations. First, we find that in average B0SMC has higher ESS than TDS consistently at different outlier levels. Second, B0SMC triggers substantially less resampling, implying that the proposal in Equation (14) is more efficient. Third, the initial ESS at $t=0$ of TDS is significantly low. This is due to TDS’ twisting approximation in Equation (7) which is highly erroneous when $T-t$ is large. This problem is also reflected in Figure 2. We see that TDS at $t=0$ has degenerate samples, and in time, the sampler essentially is moving toward a wrong path, and does not recover the true distribution at $T$. On the other hand, the marginal sequence of B0SMC smoothly moves from the reference distribution at $t=0$ and correctly hits the target distribution at $T$.
@@ Appendix A Experiment details
where we chose $a=-1$ and $b=\sqrt{2}$. We choose the auxiliary observation process to be the same as this noising process, and as a result, $A_{k}=\exp(a\,(t_{k}-t_{k-1}))$ and $\Sigma_{k}=\frac{b^{2}}{2\,a}\bigl{(}\exp(2\,a\,(t_{k}-t_{k-1}))-1\bigr{)}$. With the Gaussian mixture $\pi$, the reference distribution $p_{T}=\sum_{i=1}^{10}\gamma^{T}_{i}\,\mathrm{N}(m_{i}^{T},\Lambda_{i}^{T})$, where $m_{i}^{T}=\exp(a\,T)\,m_{i}$, and $\Lambda_{i}^{T}=E_{i}\operatorname{diag}(\theta_{i}^{T})\,E_{i}^{\mkern-1.5mu\mathsf{T}}$, where $(E_{i},\theta_{i})$ is the eigendecomposition of $\Lambda_{i}$ and $\theta_{i}^{T}=\exp(2\,a\,T)\,\theta_{i}+\frac{b^{2}}{2\,a}\bigl{(}\exp(2\,a\,T)-1\bigr{)}$. Using the same routine by replacing $T$ with $t$, the Gaussian mixture marginal $p_{t}$ is available in closed form, so does its score. We then use an Euler–Maruyama discretisation at 100 steps $0=t_{0}<t_{1}<\cdots<t_{N}=T$ to simulate the denoising process above.

### v2

@@ Abstract
Generative diffusion models, Feynman–Kac models, conditional sampling, sequential Monte Carlo,
@@ 1 Problem formulation
The overall goal of the paper is to sample from the conditional/posterior distribution $\pi(\cdot{\;|\;}y)\propto f(y{\;|\;}\cdot)\,\pi(\cdot)$, where $f$ is a given likelihood function, and $y\in\mathbb{R}^{c}$ denotes an observation (Zhao et al.,, 2025). This task is at the core interests of Bayesian statistics in general, and in the context of generative diffusion models specifically, this has enabled many downstream applications, such as image restoration problems (Luo et al.,, 2025). Recently, a few methods have been proposed to address this problem, notably the diffusion posterior sampling (Chung et al.,, 2023) and their variants (see survey by Daras et al.,, 2024). These methods rely on constructing a conditional diffusion process characterised by transition $\widetilde{q}_{k{\;|\;}k-1}(u_{k}{\;|\;}u_{k-1},y)$ that encodes the likelihood $f$ and observation, to target $\pi(\cdot{\;|\;}y)$. However, these methods are intrinsically biased due to necessary approximations to $\widetilde{q}_{k{\;|\;}k-1}$. This has then motivated another class of methods based on Feynman–Kac models and sequential Monte Carlo (SMC, Chopin and Papaspiliopoulos,, 2020), to remove such biases (e.g., Wu et al.,, 2023; Cardoso et al.,, 2024; Janati et al.,, 2024; Kelvinius et al.,, 2025; Corenflos et al.,, 2025). At the heart, they work by using $\widetilde{q}$ as a proposal, and then correct the proposal samples with importance weights. While this type of methods has shown to be empirically working well, their statistical efficiencies (e.g., effective sample size) can be significantly degraded by high dimension $d$ and informative likelihood. This challenge has been recognised in the SMC literature, such as the seminal work by Del Moral and Murray, (2015). However, in the context of SMC conditional sampling of generative diffusion models, this is under investigated. As such, the specific purpose of this paper is to address the problem when the likelihood being highly informative or when the observation is an outlier.
@@
The paper is organised as follows. In Section 2 we introduce necessary premises and explain how the conditional sampling works with a Feynman–Kac model and SMC sampler. Then, in Section 3 we show our construction of the Feynman–Kac model to facilitate sampling under outlier observations, followed by experiments in Section 4.
@@ 2 Generative Feynman–Kac models
where $M_{0}$ is the initial distribution, $Z_{N}$ is the normalising constant, and $M_{k{\;|\;}k-1}$ and $G_{k}$ are Markov (proposal) transition distribution and potential function, respectively. Given the model components, a natural approach for sampling $Q_{0:N}$ is sequential Monte Carlo (SMC) which exploits the model’s temporal structure, see Algorithm 1 for implementation. At its heart, this can be viewed as a tempering sequence, gradually evolving from $Q_{0}$ to $Q_{N}$. For detailed expositions of Feynman–Kac models and SMC samplers, we refer the readers to Chopin and Papaspiliopoulos, (2020) and Del Moral, (2004).
@@
A canonical design, perhaps the most used, of the twisting function (Wu et al.,, 2023) is given by
@@
such that the marginal distribution $Q_{k}(u_{k})=\pi(u_{k}{\;|\;}y)$. Notably, The resulting Feynman–Kac model gives an Eulerian representation of the conditional score stochastic differential equation (SDE, Song et al.,, 2021, Eq. 14) at discrete times when the denoising model $q_{k{\;|\;}k-1}$ is given by a score SDE. Precisely, if the noising process is an SDE $\mathop{}\!\mathrm{d}X_{t}=a(X_{t},t)\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}B_{t}$ with $X_{0}\sim\pi$, then $Q_{0:N}$ is the finite-dimensional distribution of the conditional SDE
$\begin{split}\mathop{}\!\mathrm{d}U_{t}&=-a(U_{t},T-t)+\operatorname{\nabla\!}\log p_{T-t}(U_{t})+\operatorname{\nabla\!}\log l^{y}_{T-t}(U_{t})\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}W_{t},\\ U_{t}&\sim\pi_{\mathrm{ref}}(\cdot{\;|\;}y),\end{split}$ (8)
[…]
However, many criticism can be said to this canonical twisting. First, in almost all practical cases, the canonical twisting has to be approximated, and the approximation errors are substantial and difficult to control particularly when NkN-k is large. Second, even if we were able to compute the canonical twisting exactly, constructing an effective proposal $M_{k{\;|\;}k-1}$, see Remark 2.1, is hard as well. These two problems are especially pronounced when the likelihood $f$ is informative, or $y$ is an outlier observation. To build some intuition, consider a heuristic: suppose that we have $J$ samples $\{U^{j}_{0}\}_{j=1}^{J}\sim q_{0}$, then, to get weighed samples for $Q_{0}(\cdot)\propto l^{y}_{0}(\cdot)\,q_{0}(\cdot)$, Equation (7) is essentially pushforwarding $\{U^{j}_{0}\}_{j=1}^{J}$ to $\{U^{j}_{N}\}_{j=1}^{J}$ and then obtain the weights by evaluating $f(y{\;|\;}\{U^{j}_{N}\}_{j=1}^{J})$ which are likely to be degenerate under outlier $y$.
@@ 3 Construction of twisting sequence with observation path
After all, if our primary interest is only the marginal $q_{N}(\cdot{\;|\;}y)$, we are not obliged to constrain $Q_{0:N}$ to follow the SDE representation in Equation (8). In this section, we depart from the canonical design, and develop a new construction of the twisting functions $\{l^{v_{k}}_{k}\}_{k=0}^{N}$, especially tailored for informative likelihood or outlier observation. Our gist consists in designing $l^{v_{k}}_{k}$ in a way that it is an “appropriate” likelihood for $q_{k}$ at $k$, by making a smooth bridging/tempering between $l_{0}^{v_{0}}$, the starting likelihood, and $l_{N}^{v_{N}}$, the target likelihood. To clarify, for instance, if $q_{0}$ is a standard Gaussian, then we ideally would like $l_{0}^{v_{0}}\approx\mathrm{N}(0;0,I)$ facilitating importance sampling at this step. We achieve this by drawing inspiration from Corenflos et al., (2025), Dou and Song, (2024), and Trippe et al., (2023) to construct an auxiliary path of observations correlated with the diffusion model. The intermediate likelihood will then reflect this correlation at all steps.
We define a new process $\{Y_{k}\}_{k=0}^{N}$ initialised with the given observation $Y_{0}=y\equiv y_{0}$, and simulate the path $\{Y_{k}\}_{k=0}^{N}$ by a noising transition $p_{k{\;|\;}k-1}(y_{k}{\;|\;}y_{k-1})$. [^1] For simplicity, we let the noising process for $y$ be the same as in Equation (1) for $x$, but they do not have to be the same in practice. Write its time-reversal as $\{V_{k}=Y_{N-k}\}_{k=0}^{N}$. We design the twisting function as a time-pushforward of the target likelihood:
@@ 3.1 Construction with Gaussian likelihood models
which is in general intractable due to the denoising $p_{0}(x_{0}{\;|\;}x_{k})$. It seems that the same problem of the canonical twisting in Equation (7) persists, as existing approximations to $p_{0}(x_{0}{\;|\;}x_{k})$ will result in large errors due to the time leap from $k$ to 0. However, it turns out that we can obtain $p_{k}(y_{k}{\;|\;}x_{k})$ based on $p_{k-1}(y_{k-1}{\;|\;}x_{k-1})$ for which we can then recursively compute $p_{k}(y_{k}{\;|\;}x_{k})$ starting at the target likelihood $p_{0}(y_{0}{\;|\;}x_{0})=f(y_{0}{\;|\;}x_{0})$. This is shown in the following proposition.
@@
If the observation noising process $p_{k-1}(y_{k}{\;|\;}y_{k-1})$ is $\phi$-stationary: $\int p_{k-1}(y_{k}{\;|\;}y_{k-1})\mathop{}\!\mathrm{d}\phi(y_{k-1})=\phi(y_{k})$, then $\mathcal{K}(\phi)(y,\cdot)=\phi(y)$ is a fixed point. The resulting twisting function will be approximately a standard Normal if the noising is an OU process.
Recall that the canonical twisting in Equation (7) requires global approximations to $p_{0}(x_{0}{\;|\;}x_{k})$ between 0 and $k$, which is hard. In contrast, the interpolating twisting in Equation (11) only needs local approximations between continuum k1k-1 and $k$, significantly reducing complexity and improving numerical robustness in practice. We show a particular case when the target likelihood is linear Gaussian for how to approximate $\mathcal{K}^{k}(f)$ in closed form.
@@
From now on, we will use the proposed twisting function
@@
in the reverse-time notation. Although the twisting is approximate, the marginal $q_{N}(\cdot{\;|\;}y)$ that we are primarily concerned is left un-approximated. Furthermore, this twisting function gives us a tractable guided proposal $M_{k{\;|\;}k-1}$ mentioned in Remark 2.1 allowing for efficiently sampling:
@@ 2 Reverse $v_{k}=y_{N-k}$ for $k=0,1,\ldots,N$
3 Compute $\{F_{k},z_{k},\Omega_{k}\}_{k=1}^{N}$ in Proposition 3.2
4 Using $M_{k{\;|\;}k-1}$ and $G_{k}$ in (14) and (15)
@@ 4 Experiments
In this section we validate the proposed B0SMC sampler in Algorithm 2 on a high-dimensional conditional sampling problem ($d=256$). Importantly, we focus on gauging how statistically robust the method is against different levels of outlier observation and number of particles. The implementation is performed on an NVIDIA A100 GPU using JAX (Bradbury et al.,, 2018), and the code is published at https://github.com/zgbkdlm/gfk for reproducibility.
The target posterior distribution is defined by a Gaussian mixture posterior
@@
For more in-depth comparison, we also plot the ESS and time-evolution of marginal $Q_{k}$ in Figures 1 and 2. The first figure gives us three observations. First, we find that in average B0SMC has higher ESS than TDS consistently at different outlier levels. Second, B0SMC triggers substantially less resampling, implying that the proposal in Equation (14) is more efficient. Third, the initial ESS at $t=0$ of TDS is significantly low. This is due to TDS’ twisting approximation in Equation (7) which is highly erroneous when TtT-t is large. This problem is also reflected in Figure 2. We see that TDS at $t=0$ has degenerate samples, and in time, the sampler essentially is moving toward a wrong path, and does not recover the true distribution at $T$. On the other hand, the marginal sequence of B0SMC smoothly moves from the reference distribution at $t=0$ and correctly hits the target distribution at $T$.
@@ Appendix A Experiment details
where we chose a=1a=-1 and $b=\sqrt{2}$. We choose the auxiliary observation process to be the same as this noising process, and as a result, $A_{k}=\exp(a\,(t_{k}-t_{k-1}))$ and $\Sigma_{k}=\frac{b^{2}}{2\,a}\bigl{(}\exp(2\,a\,(t_{k}-t_{k-1}))-1\bigr{)}$. With the Gaussian mixture $\pi$, the reference distribution $p_{T}=\sum_{i=1}^{10}\gamma^{T}_{i}\,\mathrm{N}(m_{i}^{T},\Lambda_{i}^{T})$, where $m_{i}^{T}=\exp(a\,T)\,m_{i}$, and $\Lambda_{i}^{T}=E_{i}\operatorname{diag}(\theta_{i}^{T})\,E_{i}^{\mkern-1.5mu\mathsf{T}}$, where $(E_{i},\theta_{i})$ is the eigendecomposition of $\Lambda_{i}$ and $\theta_{i}^{T}=\exp(2\,a\,T)\,\theta_{i}+\frac{b^{2}}{2\,a}\bigl{(}\exp(2\,a\,T)-1\bigr{)}$. Using the same routine by replacing $T$ with $t$, the Gaussian mixture marginal $p_{t}$ is available in closed form, so does its score. We then use an Euler–Maruyama discretisation at 100 steps $0=t_{0}<t_{1}<\cdots<t_{N}=T$ to simulate the denoising process above.